# MA-RLHF: Reinforcement Learning from Human Feedback with Macro Actions

### v1

@@ D.2 Impact of RLHF on Reward Score Distribution
We also apply Best-of-$N$ sampling on both vanilla PPO and MA-PPO. The RM score shifting patterns for these methods are illustrated in Figure 20. From the results, we can conclude that Best-of-$N$ sampling continues to enhance the performance of RLHF models effectively.
@@
D.3 Impact of Sampling Temperature
@@ Appendix E MA-RLHF Algorithms
Figure 22 illustrate the framework of MA-RLHF. In practice, to implement MA-RLHF, once the macro actions are obtained via the termination function, we compute their value (as estimated by the critic model) and rewards (based on a per-token KL penalty) using the value function estimation. With these values and rewards, we apply Generalized Advantage Estimation (GAE) without modification to derive advantage estimates and state-action value functions. These advantage estimates and state-action value functions are then used to all tokens within the macro action during the optimization of both the policy and critic models. The macro action RLHF algorithm, utilizing PPO, is detailed in Algorithm 1.
@@ F.2 Human Evaluation
In the context of HH-RLHF, annotators focus on the helpfulness of the responses. They are instructed to choose the response that provides more useful information for a given dialogue, while excluding superficial replies such as ”You’re welcome.” For the WebGPT dataset, the primary evaluation factor is factual accuracy. Annotators are provided with retrieval information relevant to the question from the dataset to aid in their judgment. They are tasked with selecting the answer that most accurately matches the retrieved information.
@@
For evaluations on the TL;DR and HH-RLHF datasets using 7B models, we conduct the human evaluation with 3 different annotators and collect their preference data to report the win rates. For all other human evaluations, we conduct them with a single annotator.

### v2

@@ D.3 Impact of RLHF on Reward Score Distribution
We apply Best-of-$N$ sampling on both vanilla PPO and MA-PPO. The RM score shifting patterns for these methods are illustrated in Figure 22 (Left and Mid Left). From the results, we can conclude that Best-of-$N$ sampling continues to enhance the performance of RLHF models effectively.
In Figure 22 (Mid Right and Right), we compare the MA-PPO with vanilla PPO using settings of $n=5$ and $n=\infty$, both of which demonstrate positive effects on the RM score distribution.
@@
D.4 Impact of Sampling Temperature
@@ Appendix E MA-RLHF Algorithms
Figure 24 illustrates the framework of MA-RLHF. In practice, to implement MA-RLHF, once the macro actions are obtained via the termination function, we compute their value (as estimated by the critic model) and rewards (based on a per-token KL penalty) using the value function estimation. With these values and rewards, we apply Generalized Advantage Estimation (GAE) without modification to derive advantage estimates and state-action value functions. These advantage estimates and state-action value functions are then used to all tokens within the macro action during the optimization of both the policy and critic models. The macro action RLHF algorithm, utilizing PPO, is detailed in Algorithm 1.
@@ F.2 Human Evaluation
For evaluation on TL;DR dataset, the annotators should first compare the overall quality of two responses. If overall qualities are equally good for responses, then they should choose the winner based on hallucination and verbosity.
In the context of HH-RLHF, annotators focus on the helpfulness of the responses:
  1. 1.
    Instruction Following: whether the generated response follows the requirements in the instruction
  2. 2.
    Usefulness: whether the advices in the response are applicable, and does the response ideally guide the user on what to do next.
Annotators are instructed to choose the response based on these aspects, while excluding superficial replies such as ”You’re welcome.” For the WebGPT dataset, the primary evaluation factor is factual accuracy. Annotators are provided with retrieval information relevant to the question from the dataset to aid in their judgment. They are tasked with selecting the answer that most accurately matches the retrieved information.
@@
For evaluations on the TL;DR and HH-RLHF datasets using 7B models, we conduct the human evaluation with 3 different annotators and collect their preference data to report the win rates. For all other human evaluations, we conduct them with a single annotator. The inter-rater agreement achieves an average of 68% on total 100 samples. On the TL;DR dataset the agreement is 64%, and on the HH-RLHF dataset the agreement is 72% across 50 samples per task.